Let (-4,-1) be a point on the terminal side of an angle theta in standard position. Draw the angle then find the 6 trig ratios as well as the measure of angle theta. Thanks.
Answer by stanbon(75887)

Let (-4,-1) be a point on the terminal side of an angle theta in standard position. Draw the angle then find the 6 trig ratios as well as the measure of angle theta.
------
x = -4
y = -1
r = sqrt[4^2 + 1^2] = sqrt(17)
------
sin(t) = y/r = -1/sqrt(17)
cos(t) = x/r = -4/sqrt(17)
tan(t) = y/x = 1/4
-----
csc(t) = -sqrt(17)
sec(t) = -sqrt(17)/4
cot(t) = 4
--------------------
Cheers,
Stan H.
--------------